OpenStudy (anonymous):

Can someone teach me how to factor trinomials clearly so i can understand. 3x^2 −42x+147

OpenStudy (nubeer):

hmm yes a few ways.. maybe you can post question so somone can help on that and then you can understand how to do it.

OpenStudy (nubeer):

are you familiar with quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (nubeer):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{2a}\] this is the formula for finding roots or factors

OpenStudy (nubeer):

this formula is only applied when the highest power of x is 2 like in this question of yours.

OpenStudy (nubeer):

3x^2 −42x+147 this is your question a = 3, b = -42 , c = 147 just plug this in formula and u will get an answer.

OpenStudy (nubeer):

getting it? @monica123